the equation of a line with the slope 1.3 that passes through the point (-2,0)
Answer by rfadrogane(214)

the equation of a line with the slope 1.3 that passes through the point (-2,0)
sol'n:
by using point-slope form,
y-y1 = m(x-x1)@ P(-2,0), m = 1.3 or 13/10
y-0 = 13/10(x+2)
10y = 13(x+2)---answer (point-slope form)